Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 25

Observer: JHill (LBTO), DMiller (Tucson), plus JStorm (Potsdam)
Telescope Operator: WWack
Instrument Support: DOfficer
Telescope Support:
SW support: (Tucson)

Summary

Things were pretty good at the beginning, but it only went downhill from there.

AGW transform is close to correct, but there is a problem that the star position is Position Angle dependent. We suspect a PCS problem because we can't see any way that oacontrol knows the Position Angle. BUT, the problem that was pronounced on StoneA seems to have dissappeared on the 12 points we got on StoneC. ???

JMH: That may imply a coordinate problem with the central star of StoneA, although it is the central star of StoneC with a unknown proper motion. We discovered the next day that the known proper motion for StoneA was missing from the LBTO copy of the catalog which we were using for the first part of this night. The 0.6 arcsec shift in the position seems not quite large enough to explain the effect we saw, but it may be a contributor.

Plenty of problems with SX Primary mirror panicing.

Some minor problems with GCSL and GCSGUI.

And we closed a couple times for weather.

Details

All Times in UT

00:30 Open vent doors

Open

00:50 Wayne has completed the closed dome test of telescope motions and the chamber is now open

00:54 After two attempts of sending a TRACK preset (GCSL not running and Rotator not Ready) a successful preset is sent. Lost 5 minutes to these telescope not prepared issues.

00:55 pointto M5_1526 TelMode=TRACK gs=0

00:57 ptspiral executed by Wayne to find star on IRTC. Oops, presetTelescope not in PARALLACTIC mode

00:58 pointto M5_1526 TelMode=TRACK gs=0 PAM=no

00:59 ptspiral executed by Wayne to find star on IRTC. Found and centered, IE=7 CA=17. A lot of coma!

01:02 pointto M5_1526 TelMode=ACTIVE gs=2 PAM=yes. Guide star less than 2" from hotspot. Guiding and Active running. Several thousand nm of coma. Down to RMS of ~500 in about 5 iterations

01:10 pointto M5_1526 TelMode=ACTIVE gs=2 PA=220 (on y-axis was 268) to see if AGw transform is close to correct. Star <1" from hotspot. Collimate for 5 minutes while John gets pointing model data collection routine setup.

01:17 Wayne turned on pointing log: /home/telescope/Pointing_250ct09

01:20 John tries to take pointing model data. Gets a seg-fault from his gopointing script in TRACK mode. Not clear why. (Later we worked out that Pointingstars5 returns 999 when it has no star in that grid. This makes the preset fail. Haven't yet explained the seg fault.)

01:30 pointto first pointing star at 25 El manually. Star not found. John performing a ptspiral. Found and centered with IE=-13 CA=-43

01:38 Took first pointing model point (0206) manually

UT StarID IRTC Cube IRTC SIngle Comment
1:38 act0206 0009 013804  
1:42 act0407 0010 014214  
1:45 wt10_401 0011 014526  
1:50 wt10_332 0012 014959 Binary Star?
1:53 act0281 0013 015336 Looked oblong
2:00 wt10_196 0014 020030 coma and focus, collimate on next
Collimated and set IE=1.0 CA=10.0
2:15 Dark 0015 201552 IRTC image with probe in
2:17 act253 0016 021738  
2:24 wt10_253 0017 022414  
2:31 act0083 0018 023131 Barely on Acq. Image; OK
2:35 wt10_151 0019 023526  
2:40 act0247 0020 024014  
2:44 act0208 0021 024418 No Star
2:47 wt10_340 0022 024742  
2:50 wt10_325 0023 025031  
2:58 act0208 0024 0257  
3:00 act0261 0025 030029  
3:03 act0291 0026 030324  
3:12 act0291 0027 031209  

02:06 manual pointto act0253 TelMode=ACTIVE gs=0. Did not acquire the correct star: too dim. Collimate a few times to improve image.

02:12 pointto in ACQUIRE mode to find the correct star. Set IE=1.0 CA=10.0

02:15 Take IRTC set with probe on-axis. Can be used as a dark.

02:17 new pointing model data point

02:27 gopoint act0083 Star not found. Set IE=-7 CA=-20

02:30 gopoint act0083 Star on edge of acquisition image, and pulled to center. Data collected (see above)

02:53 pointto act0208 ACQUIRE. Set IE=-10. CA=-45

02:58 gopoint act0208. Acquired Pointing data. Set WFSC Exp time to 10000 ms on the command line for these bright stars

03:01 set WFSExpTime to 10000 ms in /lbt/tcs/current/etc/GCS/AIP_IRTC_DG_L.cfg

3:02 Primary Mirror Panic. Too much force. act0261 images are blank. We think there was -300 Z22 left on the primary from last night with LBC. All of our Active Optics corrections have been small amounts of Z11.

03:24 pointto act0261 TelMode=ACQUIRE EL=77 IE=-7 CA=28

03:25 pointto M5_0855 TelMode=ACTIVE gs=0 EL=67 Star not found

03:26 pointto M5_0855 TelMode=ACQUIRE gs=0 Wayne set IE=4.0 CA=17.0

03:28 pointto M5_0855 TelMode=ACTIVE gs=0

03:29 Make IssueTrak /Repository/20091025/GCS/acquire000022.fits Star out of focus and found in acquisition image. However, error reported by GCS when trying to star guiding at 3:29 "Could not adjust exposure time" guider images /Repository/20091025/GCS/guiderimage000847-000853.fits

AGW Calibration on StoneA

03:43 pointto M5_0036 TelMode=ACQUIRE (near StoneA) IE=0 CA=-10 EL=53

03:45 pointto StoneAagw TelMode=ACQUIRE IE=2.0 CA=-19.0 EL=44

03:46 pointto StoneAagw TelMode=ACTIVE gs=12 Collimating

Stopped the pointing log.

Seeing has degraded back to 1 arcsec at EL=46 degrees. Taking 1 sec x 45 cubes. Seeing got as bad as 1.5 arcsec during this set of data.

03:51 cl < home$tran_StoneA_00.cl
transtar (12) irtc.20091025.035241.fits cube 28
transtar (16) irtc.20091025.035459.fits cube 29
transtar (17) irtc.20091025.035726.fits cube 30
transtar (18) irtc.20091025.035948.fits cube 31
transtar (14) irtc.20091025.040224.fits cube 32
transtar (12) irtc.20091025.040447.fits cube 33
transtar (6) irtc.20091025.040711.fits cube 34
transtar (5)
transtar (3)
transtar (12)
*CA=1*
transtar (11)
CA=-10

Seeing has improved to 0.8 arcsec at EL=49. Thin clouds are rolling through.

04:17 cl < home$tran_StoneA_90.cl
transtar (12)
transtar (13)
transtar (15) irtc.20091025.042400.fits cube 41
transtar (19)
transtar (20)
transtar (21) irtc.20091025.043058.fits cube 44
transtar (16)
CA=-5
transtar (17)
transtar (18)
transtar (14) irtc.20091025.044043.fits
transtar (12)

04:44 cl < home$tran_StoneA_180.cl
transtar (13) irtc.20091025.044556.fits
transtar (15)
transtar (19)
transtar (20)
transtar (21) irtc.20091025.045502.fits
transtar (9)
transtar (2)
transtar (13) irtc.20091025.050220.fits

The problem here is that the location of the image on the IRTC is a strong function of the position angle. Within the seeing disk, the images are all aligned at a particular position angle, but they seem to trace out a 3 arcsec diameter circle as we rotate the position angle. Is this a bug in PCS? We can't see how the AGW transform could be making this error, since it doesn't know what the position angle is, and the AGW and IRTC are locked to the same physical coordinate system.

Sky is clear again, but seeing still around 1 arcsec.

05:05 offset 20 20; takeseq 1.0 Nimages=45 - irtc.20091025.050505.fits cube 58 - This is the sky background frame without any star.

05:06 cl < home$tran_StoneA_270.cl
transtar (1) irtc.20091025.050824.fits cube 59
CA=2
transtar (2)
CA=5
transtar (8)
IE=0
transtar (4)
transtar (6) irtc.20091025.051815.fits
IE=6
transtar (1)
IE=8
transtar (2) irtc.20091025.052300.fits cube 65

Repeat the PA=0 data to break the degeneracy between a postion angle problem versus a rotator angle problem. The seeing also seems to be improving to 0.6 arcsec on the guider.

05:24 cl < home$tran_StoneA_00.cl
transtar (12) irtc.20091025.052534.fits cube 66
transtar (16)
transtar (17)
transtar (18)
transtar (14)
transtar (12)
transtar (6)
transtar (5)
transtar (3)
transtar (12) irtc.20091025.054656.fits cube 75
transtar (11)

Seeing varying between 0.7 and 1.0 arcsec on guider. Sky is becoming mostly cloudy.

05:50 cl < home$tran_StoneA_90.cl
transtar (12) Star lost to clouds.
transtar (13) Star lost to clouds.

Close

05:56 Close for clouds.

John is doing closed-dome testing of the gopoint.cl script.

06:11 Both John and Doug have their GCSGUIs core dump.

Open

06:30 Reopen to a clearing sky. T=3.9 degC.

06:33 pointto M5_0036 TelMode=ACQUIRE (near StoneA) - Adjust *IE=8 CA=15 *

06:38 pointto AtoneAagw gs=12 ACTIVE PA=0 - CA=11

Seeing is fluctuating between 0.7 and 1 arcsec on guider. Sky is mostly clear with thin cirrus.

06:42 cl < home$tran_StoneA_90.cl
transtar (12) irtc.20091025.064349.fits cube 82
CA=6
transtar (13)
transtar (15)
transtar (19)
script crashes while John is editting it. Jesper joins us by phone.
transtar (20)
transtar (21) irtc.20091025.065919.fits cube cube 87 What happened to the star?
transtar (16) Couldn't move probe.
Resmsg[3]  = couldn't move probe to position (SFP) 27.987, 97.8186, focus: 28.2874 :AGw command status: timed out waiting for probe to reach destination
07:03 stop GCSL, stopAGW, startAGW

07:07 transtar (21) irtc.20091025.070843.fits cube 89 - no star.
 Resmsg[3]  = couldn't move probe to position (SFP) -115.241, 146.225, focus: 16.4663 :AGw command status: AGW4: Motor already in motion

07:08 home the probe again.

07:10 transtar (21) same AGw error repeats BUT Wayne homed unit 3 and not unit 4!

07:12 stop / start / home unit 4 this time. Stop and start GCSL.

07:17 transtar (21) - irtc.20091025.071759.fits cube 91 - SX Primary Mirror panic - no star.
Sun Oct 25 07:17:42.484 2009  primary mirror adjustMirrorCollimation failed with argument 0.0000000e+00 because left PMC Mirror GoToPositionAbs Command Failed because Wrong mirror state

Lost 40 minutes to messing with a cranky primary mirror. Not sure that we ever identified what was actually wrong, but we went through the phase of recovering HP1 from a travel out limit. And we saw the mirror apparently still chasing after a panic.

07:42 transtar (21) irtc.20091025.074333.fits cube 92
transtar (16) irtc.20091025.074642.fits cube 93

StoneC

07:48 pointo M5_0151 gs=0 ACQUIRE PAM- (landed within 1 arcsec of target! - no adjustment needed)

07:51 pointto StoneCagw gs=18 ACTIVE PA=INDEF=21.78
What does it mean when GCS reports -1 for the WFS rms? There are clouds, but the WFS still has images on it.

07:59 Stop and restart GCSL.

08:00 pointto StoneCagw gs=18 ACTIVE PA=INDEF=21.78
Finally found the GCS log message which is complaining that we have only 60 max peak counts when it wants 100.

08:07 Lowered countThreshold from 100 to 50 in AIP_IRTC_DG_L.cfg, but part of the problem is that we had turned the default integration time down to 5000 msec.

08:11 setWFSExpTime 30000 from command line, and then it sent corrections. But why wasn't the exposure time autoadjusting. Was it because we had more than 1000 counts?

08:15 Changed WFScam_lowCounts from 1000 to 2000.

08:16 cl < home$tran_StoneC_00.cl
transtar (10) irtc.20091025.081735.fit cube 94
transtar (14) Didn't get IRTC data because John was playing with video button. Clouds are making the star pretty faint on IRTC.
transtar (15)
Changed WFS default exposure from 5000 to 10000 msec.
transtar (12) irtc.20091025.082330.fits cube 96 -- CLOUDS
transtar (18) irtc.20091025.082545.fits cube 97 -- CLOUDS

08:27 Stop and restart GCSL to get the newest parameters. It doesn't seem to be picking up my changes on the fly. The default WFS exposure and countThreshold only updated when I restarted GCS.

08:28 transtar (10)

Why didn't GCS increase the exposure to the maximum in this example?
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: slopeAverage: pupil grid shifting moved the center to x = 402.95, y = 262.95 
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: 4 corner's mean background is : 1186.00
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: slopeAverage: final xCenter = 402.82, yCenter = 264.34
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: maxPeak: 1443, bkgrd: 1186
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: maxPeak : 257   avg. bkgrd: 1186   countThreshold: 50
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: couldn't adjust exposure time to 778210 ms as the maximum is defined as 30000.
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: trying to estimate counts for maximum exposure time 30000 ms.
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: estimated counts would be 771.
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: estimated counts would exceed limits. Unable to adjust exposure time.
Oct 25 08:30:11 lbtmu304 LBT_left_GCS: WFS: WFS image analyzed.

08:33 transtar (14)

08:36 PSF complaining about bad new influence forces.

08:37 Sent manual +100 nm Z11, but that was also rejected. (opposite sign of what GCS wanted)

08:38 Clear C00 for SX

08:41 cl < home$tran_StoneC_90.cl (working through clouds)
transtar (21) irtc.20091025.084316.fits cube 100
transtar (20)

Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: slopeAverage: final xCenter = 376.81, yCenter = 276.91
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: maxPeak: 1244, bkgrd: 1187
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: maxPeak : 57   avg. bkgrd: 1187   countThreshold: 50
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: couldn't adjust exposure time to 3508772 ms as the maximum is defined as 30000.
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: trying to estimate counts for maximum exposure time 30000 ms.
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: estimated counts would be 171.
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: estimated counts would exceed limits. Unable to adjust exposure time.
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: WFS image analyzed.
Oct 25 08:48:08 lbtmu304 LBT_left_GCS: WFS: calculated Zernikes for /Repository/GCSfiles/dailyimages/wfscimage000414.fits are

This became Issue 2409.

08:49 cl < home$tran_StoneC_180.cl
transtar (13)
transtar (16)
transtar (17)
transtar (7) irtc.20091025.085743.fits cube 105

Pointing

08:59 restart the pointing log Pointing2-25oct09

/OldRepository/IRTC_Data/091025_Pointing

09:00 gopoint 1 active+ - it didn't accept my preset, and then this gives IRAF the seg fault problem. The problem seems to be that gopoint is getting an invalid target name back from Pointingstars5.

IRTC> gopoint 1 active+
GOPOINT: Saving target names in: gopoint.log
GOPOINT: Presetting telescope to grid position 1 
Warning: Attempt to delete a nonexistent file (junk_gopoint.tmp)
GOPOINT: Grid: 999 05
GOPOINT: pointto TelMode=ACTIVE 05 gs=0 PA=0
   to examine the log: page LOGs/pointto_090044.log
GOPOINT: Sleeping for 30.  seconds to allow WFS time.

09:06 gopoint 1 WT10_379

Oct 25 09:10:01 lbtmu304 LBT_left_GCS: gcs.interface.azcam.wfsc.readout.failed: interrupt received while waiting for WFS cam image

09:12 Set WFSminExposure to 1000 , stop and restart GCSL to read parameters.

09:13 gopoint 1 irtc.20091025.091451.fits cube 107
The WFS exposure time adjustment did work correctly going down.

09:15 gopoint 2 ploop+ missed star at 20 deg EL

CA=-30

09:18 gopoint 2 ploop-

09:20 ACT0406 ACQUIRE

09:21 ptspiral - Adjust *IE=-8 CA=-50 *

09:25 gopoint 2 ploop- ACT0406

09:27 gopoint 3 WT10_405

09:30 retake a manual sequence 1 x 10

09:31 manual seqeunce 0.1 x 100 irtc.20091025.093116.fits

gopoint 4 = grid 999

09:33 gopoint 5 WT10_291

Fixed gopoint.cl so it recognizes that grid 999 means no star is available.

09:38 gopoint 6 WT10_172 (This is a binary star. Coords will be poor.) irtc.20091025.094422.fits

09:44 gopoint 8 WT10_091

09:49 gopoint 9 WT10_144 irtc.20091025.095118.fits

09:52 gopoint 10 WT10_284 guess CA=17 - SX Mirror panics, no data

09:59 Clear C00 after the second panic.

10:03 gopoint 7 WT10_265

10:07 gopoint 11 ACT0201 guess CA=-50 irtc.20091025.101054.fits cube 116

10:11 gopoint 12 ACT0335 - guider got a faint star near the bright one I think.

10:15 pointto ACT0335 ACQUIRE

10:11 gopoint 12 ACT0335 with guess CA=-30 irtc.20091025.101913.fits cube 118

Close

10:20 Close for impending clouds with radar echoes.

PCS closed-dome tests

This is EL=56 degrees.

10:43 pointto M5_0258 gs=0 TelM="ACQUIRE" PA=0
GS is set to SFP(x,y): 0.0842008, 0.170868

10:45 pointto M5_0258 gs=0 TelM="ACQUIRE" PA=90
GS is set to SFP(x,y): 0.170357, -0.0812043

10:47 pointto M5_0258 gs=0 TelM="ACQUIRE" PA=180
GS is set to SFP(x,y): -0.0789452, -0.170049

10:49 pointto M5_0258 gs=0 TelM="ACQUIRE" PA=270
GS is set to SFP(x,y): -0.169591, 0.0762538

At face value, this seems to produce a 0.4 mm diameter circle. But we need to be careful about wavelength. PCS is assuming 0.63 versus 1.4 microns, while we are observing tonight with 0.63 (r') versus 1.2 microns (J).

Now switching to EL=39 degrees.

10:53 pointto M5_0243 gs=0 TelM="ACQUIRE" PA=270
GS is set to SFP(x,y): -0.343213, 0.0538687

10:56 pointto M5_0243 gs=0 TelM="ACQUIRE" PA=90
GS is set to SFP(x,y): 0.342903, -0.05064

Now, if the sign of this refraction correction were reversed, then we might have a circle that is big enough to be the explanation.Positive PCS_Y is North (higher EL) at PA=0, so it seems correct that the guide probe is also positioned at higher EL to see the bluer guide star compared to the redder IRTC image.

11:18 pointto M5_0243 gs=0 TelM="ACQUIRE" PAM-
GS is set to SFP(x,y): -1.90255e-09, 0.340712

-- DougMiller - 21 Oct 2009
I Attachment Action Size Date Who Comment
daq.20091025.gifgif daq.20091025.gif manage 42 K 26 Oct 2009 - 00:48 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r11 - 12 Nov 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback